OpenStudy (anonymous):

what is 6x/45 in simplest form?

OpenStudy (anonymous):

Thanks!!!!

OpenStudy (anonymous):

Sorry, that was wrong, let me just redo it

OpenStudy (anonymous):

\[\frac{ 6x }{ 45 } = \frac{ 3(2x) }{ 3*15 } = \frac{ 2x }{ 15 }\]